# equal contribution

### v1

@@ Appendix 0.A Video Results
We include an extensive set of 4D generation results in video format in the supplementary website, best viewed with Google Chrome. There, we showcase compositional 4D scenes, a diverse set of trajectory-conditioned 4D results, comparisons to 4D-fy [4] and DreamGaussian4D [71], as well as ablation studies motivating our design choices.
@@ Appendix 0.E Geometry Visualization
We show that TC4D generates scenes with comparable geometry to 4D-fy while introducing significantly more motion in Figs. 10 and  11. Note that the focus of this work is not to improve or provide high-quality geometry but introduce more realistic and larger-scale motion.

### v3

@@ Appendix 0.A Video Results
We include an extensive set of 4D generation results in video format in the supplementary website https://sherwinbahmani.github.io/tc4d, best viewed with Google Chrome. There, we showcase compositional 4D scenes, a diverse set of trajectory-conditioned 4D results and comparisons to 4D-fy [4].
@@ Appendix 0.B Implementation Details
Camera pose sampling.
For training, we sample one random view per VSDS update shared across all trajectory timesteps. For each iteration, we randomly sample a camera on a sphere with radius in [1.5, 2.0], field of view in [15, 60], elevation in [0, 30], and azimuth in [-180, 180]. For inference, we render videos with radius 1.75, field of view 40, elevation 15, and equidistant azimuth samples in [-180, 180]. The radius and field of view are selected to ensure that objects largely stay within the image frames.
@@ Appendix 0.D Automated Generation
Moreover, we show single-object automated results in Fig. 10. Improving automated trajectory generation and multi-object text-to-4D generation using, e.g., scene graphs [22], is an important direction for future work.
@@
Figure 10: Results using automated trajectory generation.
@@ Appendix 0.E Geometry Visualization
We show that TC4D generates scenes with comparable geometry to 4D-fy while introducing significantly more motion in Figs. 11 and  12. Note that the focus of this work is not to improve or provide high-quality geometry but introduce more realistic and larger-scale motion.